OpenStudy (calculusxy):

A scale on a map is 1 1/2 inches = 10 miles. How many miles does a 21 inch line represent?

OpenStudy (calculusxy):

Here are the options: a. 31 1/2 b. 140 c. 0.714 d. 150 e. none of the above *I chose A. 31 1/2.

OpenStudy (anonymous):

11/2 for 10 miles 21 for 10/(11/2)*21=420/11 miles

OpenStudy (calculusxy):

i meant one and a half not eleven over two

OpenStudy (anonymous):

oh..then 10/(3/2)*21=70

OpenStudy (calculusxy):

ok so i here is another question: A 10-minute call to Hawaii after 9 p.m. costs $5.75. if the charges for the first 3 minutes was $3.65, what was the charge for each additional minute?

Directrix (directrix):

A scale on a map is 1 1/2 inches = 10 miles. How many miles does a 21 inch line represent? (3/2)/10 = 21/x which means 1.5 inches is to 10 miles as 21 inches is to x miles. (3/2)/10 = 21/x @apurvasijaria I cranked this out and did not get 70 as you reported above >> oh..then 10/(3/2)*21=70 Would you show me the error in my work? Thanks.

Directrix (directrix):

@calculusxy >>> *I chose A. 31 1/2. Not correct

OpenStudy (anonymous):

@Directrix I cranked this out and did not get 70 as you reported above sorry my mistake its 140.

Directrix (directrix):

@apurvasijaria Thanks for checking.